OpenStudy (aaalive):

Express your answer in lowest terms 1/6 x 5/7=?

pooja195 (pooja195):

\[\huge~\rm~\frac{ 1 }{ 6 }\times \frac{ 5 }{ 7 }=\frac{ (5)(1) }{ (6)(7)}=\frac{ ? }{ ? } \]

OpenStudy (aaalive):

I still don't get it....sorry! can you let me know what the parentheses are for?

OpenStudy (yaya090600):

the parentheses mean you're multiplying

OpenStudy (aaalive):

okay, thanks!

OpenStudy (yaya090600):

no problem

pooja195 (pooja195):

Sorry if i was confusing you!

OpenStudy (aaalive):

let me try that And i will come back with what I think it is...

OpenStudy (aaalive):

is it 5/42?

OpenStudy (yaya090600):

yes

OpenStudy (aaalive):

yay! thanks, I gave you a medal!

OpenStudy (yaya090600):

you're welcome :)

OpenStudy (bamboo888):

5/42

OpenStudy (bamboo888):

5/7 x 1/6 = 5/42

OpenStudy (aaalive):

thank you, @Bamboo888 but I already got help... but thanks anyways!